# ”Security for Everyone” in Finite Blocklength IRS-aided Systems With Perfect and Imperfect CSI

### v2

@@ IV-3 Complexity Analysis
$\mathcal{O}_{\boldsymbol{w}}$ $\triangleq\mathcal{O}\left\{\sqrt{{b_{1}}+2}({c})({c}^{2}+{c}{b_{2}}+{b_{3}}+({c}+1)^{2})\right\},$ (64)
$\mathcal{O}_{\boldsymbol{\Phi}}$ $\triangleq\mathcal{O}\left\{\sqrt{{b_{1}}+4N}(2N)(4N^{2}+2N{b_{2}}+{b_{3}}+4NM)\right\},$ (65)
@@
${b_{1}}$ $\triangleq{\sum}_{k}^{\mathcal{K}}({a_{1}}+{a_{2}})+2k({a_{2}}+{a_{3}})+(k-2)({a_{1}}+{a_{3}}),$
${b_{2}}$ $\triangleq{\sum}_{k}^{\mathcal{K}}({a_{1}}^{2}+{a_{2}}^{2})+2k({a_{2}}^{2}+{a_{3}}^{2})+(k-2)({a_{1}}^{2}+{a_{3}}^{2}),$
${b_{3}}$ $\triangleq{\sum}_{k}^{\mathcal{K}}({a_{1}}^{3}+{a_{2}}^{3})+2k({a_{2}}^{3}+{a_{3}}^{3})+(k-2)({a_{1}}^{3}+{a_{3}}^{3}).$
@@ V Numerical Results
The convergence tolerance $\epsilon_{t}$ $10^{-3}$
Sim. initial settings $\beta_{k}^{(\iota)}=1$, $\beta_{ke}^{(\iota)}=1$, $o^{(1)}=10$, $o_{\text{max}}=30$, $\xi_{k}=\xi_{e}=0.02$.
@@
In this section, we perform an extensive simulation to evaluate the performance of the proposed approach. The results are obtained using the MATLAB and CVX toolboxes. In this setup, Alice is located at $(15,0,15)$, the IRS is located at $(0,25,40)$, and the users are randomly distributed to the right of Alice over a $(60m\times 60m)$ area. The Eve (e.g., a compromised legitimate user) is randomly located in $(100m\times 100m)$ outside the users’ area. Note that if Eve is too close to one of the users, it is mainly impossible to guarantee the positive SR for all the users [38]. In this case, other methods, e.g., encryption or friendly jamming, can provide users’ secrecy [14].
@@
The Alice-to-IRS direct path loss factor is ${\beta_{{}_{\text{{AR}}}}}\triangleq\boldsymbol{G}_{\text{\text{A}}}+\boldsymbol{G}_{\text{IRS}}-35.9-22\log_{10}(d_{{}_{\text{{AR}}}})$ in dB, where $d_{{}_{\text{{AR}}}}$ is the distance between Alice and the IRS in meters, while $\boldsymbol{G}_{\text{\text{A}}}$ is Alice antenna gain, and $\boldsymbol{G}_{\text{IRS}}$ is the IRS elements’ antenna gain [39]. The path loss factor from the IRS to the users and the Eve is ${\beta_{\text{R}i}}\triangleq\boldsymbol{G}_{\text{IRS}}-33.05-30\log_{10}(d_{\text{R}i})$ dB, for $i\in\{k,e\}$, $d_{\text{R}i}$ is the distance between the IRS and the users and Eve in meters. The spatial correlation matrix is $[\boldsymbol{R}_{Ri}]_{l,\bar{l}}\triangleq\exp{(j\pi(l-\bar{l})\sin\hat{\vartheta}\sin\hat{\aleph})}$ for $i\in\{k,e\}$, where $\hat{\aleph}$ is the elevation angle and $\hat{\vartheta}$ is the azimuth angle [19]. The elements of the Alice-to-IRS channel are generated by $[\boldsymbol{G}_{{}_{\text{{AR}}}}]_{a,b}\triangleq\exp({j\pi\left((b-1)\sin\overline{\Theta}_{b}\sin\overline{\vartheta}_{b}+(a-1)\sin(\Theta_{n})\sin(\vartheta_{b})\right)})$, where ${\Theta}_{n}\in(0,2\pi)$, ${\vartheta}_{n}\in(0,2\pi)$, and $\overline{\Theta}_{n}\triangleq\pi-\Theta_{n}$, and $\overline{\vartheta}_{n}\triangleq\pi+\vartheta_{n}$ [19]. The small-scale fading channel gain $\widehat{\boldsymbol{u}}_{i}$ for $i\in\{k,e\}$ is modeled as a Rician fading channel with K-factor$=3$.
@@
We define the CSI error bounds as $\Omega_{i}\triangleq\delta_{i}\|\widehat{\boldsymbol{u}}_{i}\|_{2},\forall k$, where $\delta_{i}\in[0,1),i\in\{k,e\}$ is the relative amount of CSI uncertainty. When $\delta=0$, Alice can obtain the perfect CSI of the IRS to users/Eve reflected channel, the transmission duration is $t_{t}=0.1$ms which is suitable for FBR transmission [40], and the choice of $0.1$ ms end-to-end delay ensures having a quasi-static channel during FBR communication [41], $\xi_{i}=10^{(-5)}$, the bandwidth $B=1$Mhz. Unless stated otherwise, the simulation parameters are defined in Table I. We multiply the results by $\log_{2}(e)$ to convert them to bps/Hz. Lastly, for comparison purposes, we compare our Max-Min algorithm against the SSR maximization algorithms. Due to the page limitation, the steps to maximize SSR has been omitted, however, the SSR maximization can be tackled by setting the objective function in problem ($\mathcal{P}1$) and problem ($\mathcal{P}2$) to $\underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\sum_{k}^{\mathcal{K}}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta})$.
@@
Figure 2: Convergence rate of the Max-Min algorithm under prefect CSI with $M=10$, $\mathcal{K}=6$, $N=16$.
@@
Fig. 3 plots the users’ SR distribution with $M=10$, $\mathcal{K}=6$, and $N=16$ under the Max-Min algorithms. The Max-Min algorithms achieve almost the same SR for all the users when the IRS’s PREs are optimized. However, when the IRS’s PREs are not practically optimized, the algorithms fail to achieve secrecy fairness among the users. The results demonstrate the importance of properly optimizing the IRS’s PREs. The proposed algorithm for the FBR case achieves secrecy and SR fairness among the users. The results demonstrate the robustness of the proposed algorithms to achieve secrecy for all the users, even under the FBR constraints, namely, the transmission duration and the latency.
Figure 3: Users’ SR with $M=10$, $\mathcal{K}=6$, $N=16$.
Figure 4: Users’ SR with with $M=10,\mathcal{K}=6,N=16$.
Fig. 4 portrays the users’ SR under different values of the IRS to the users/Eve imperfect CSI. The Max-Min algorithm achieves secrecy fairness in the LBR/FBR cases, even when Eve’s reflected channel has higher uncertainty than the users’ reflected channel.
Figure 5: Jain’s Index Vs. the uncertainty level $\delta_{k}=\delta_{e}$ with $M=10,\mathcal{K}=5,N=16$.
[…]
@@
Figure 6: Minimum SR Vs. the uncertainty level $\delta_{k}=\delta_{e}$ with $M=10,\mathcal{K}=5,N=16$.
Figure 7: Minimum SR Vs. the number of users $\mathcal{K}$ with $M=10,N=16$, $\delta_{k}=\delta_{e}=0.02$.
@@ VI Conclusions
In this paper, we proposed a framework to achieve secure communications for all the users under the FBR constraints for IoT setting in URLLC/mMTC applications aided by IRS. Specifically, through linearization and different non-convex optimization techniques, we designed computationally efficient algorithm to maximize the minimum SR among all users under both perfect and imperfect CSI from IRS to the users and the eavesdropper. Extensive simulations results showed that the Max-Min algorithm can provide secure communications for all the users under FBR constraints even with only imperfect CSI. Note that if Eve is too close to one user, it is infeasible to guarantee secrecy for all the users and other solutions like cryptography-based or friendly jamming-based methods should be in place. In the future, one can consider a multi-hop scenario with the joint coding over multiple IRSs, similar to the cooperative MIMO setting [43].
@@ Appendix A Proof of theorem 1
First we prove that the sequence $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})$ is non decreasing for all $k$, i.e. $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})\geq\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})$ for all $\iota>0$. To that end, with the aid of the CVX solver, we obtain $\boldsymbol{w}^{(\iota+1)}$ by solving problem ($\mathcal{P}1.3$). CVX solver is guaranteed to find the optimal solution. Thus, we have $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta})>\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta})$ for all $k$. Next, from ($\mathcal{P}2$), we obtain $\boldsymbol{\theta}^{(\iota+1)}$ by solving (33), where we have $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota+1)})>\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)})$ as stated before.
@@
$\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})>\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)}),$ (66)
the optimal sequence $\{(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})\}$ thus converge to a point $\{({\boldsymbol{w}}^{*},{\boldsymbol{\theta}}^{*})\}$ which is the solution obtained from solving ($\mathcal{P}1.3$) and ($\mathcal{P}1.4$).
Next, we prove that the converged point $\hat{X}^{*}\triangleq\{\boldsymbol{w}^{*},\boldsymbol{\theta}^{*}\}$ is a locally optimal solution of problem $(\mathcal{P}1)$. For that, we show that the converged point satisfies the Karush-Kuhn-Tucker (KKT) condition of the problem. The KKT condition for problem $(\mathcal{P}1.3)$ is satisfied at $\boldsymbol{\theta}^{*}$. Let $\boldsymbol{Y}(\boldsymbol{\theta})$ be the objective function of $(\mathcal{P}2)$ and $\boldsymbol{T}({\boldsymbol{X}})=[\boldsymbol{T}_{1}({\boldsymbol{X}}),\boldsymbol{T}_{2}({\boldsymbol{X}}),\dots,\boldsymbol{T}_{I}({\boldsymbol{X}})]$ be the set of constraint of problem $(\mathcal{P}2)$. Then, we can write:
@@
$z_{i}\geq 0,t_{i}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,\forall i.$
@@
$z_{i}\geq 0,t_{i}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,\forall i.$
@@
$t_{i}\geq 0,t_{i}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,\forall i,$
@@ Appendix B Proof of Lemma 1
Let $q$ be a scalar complex variable, and $q^{(\iota)}$ is the fixed point obtained at iteration $(\iota)$, then the following inequality holds [20]:
$|q|^{2}\geq q^{*,(\iota)}q+q^{*}q^{(\iota)}-q^{*,(\iota)}q^{(\iota)}.$ (70)
By replacing $q$ with $({\boldsymbol{u}}_{k}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}$ we obtain (39). Thus, this concludes the proof. $\blacksquare$
@@ Appendix C Proof of theorem 2
Similar to the proof of Theorem 1, it can be shown the sequence $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})$ is non-decreasing for all $k$, i.e. $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})\geq\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})$ for all $\iota>0$, and the converged point is a locally optimal solution for problems. $\blacksquare$.

### v3

@@ IV-3 Complexity Analysis
$\mathcal{O}_{\boldsymbol{w}}$ $\triangleq\mathcal{O}\left\{\sqrt{{c_{1}}+2}(a)(a^{2}+a{c_{2}}+{c_{3}}+(a+1)^{2})\right\},$
$\mathcal{O}_{\boldsymbol{\Phi}}$ $\triangleq\mathcal{O}\left\{\sqrt{{c_{1}}+4M}(2M)(4M^{2}+2M{c_{2}}+{c_{3}}+4MN)\right\},$
@@
${c_{1}}$ $\triangleq{\sum}_{k}^{\mathcal{K}}b_{1}(k-1)+b_{2}(2k+1)+b_{3}(3k-2),$
${c_{2}}$ $\triangleq{\sum}_{k}^{\mathcal{K}}b_{1}^{2}(k-1)+b_{2}^{2}(2k+1)+b_{3}^{2}(3k-2),$
${c_{3}}$ $\triangleq{\sum}_{k}^{\mathcal{K}}b_{1}^{3}(k-1)+b_{2}^{3}(2k+1)+b_{3}^{3}(3k-2),$
@@ V Numerical Results
The convergence tolerances $\epsilon_{t},\epsilon_{t_{1}},\epsilon_{t_{1}}$ $10^{-3}$
Sim. initial settings $\alpha_{k}^{(n)}=1$, $\alpha_{ke}^{(n)}=1$, $a^{(1)}=10$, $a_{\text{max}}=30$, $\delta_{k}=\delta_{e}=0.02$.
@@
In this section, we perform an extensive simulation to evaluate the performance of the proposed approach. The results are obtained using MATLAB and CVX toolboxes. In this setup, Alice is located at $(15,0,15)$, the IRS is located at $(0,25,40)$, and the users are randomly distributed to the right of Alice over a $(60m\times 60m)$ area. The Eve (e.g., a compromised legitimate user) is randomly located in $(100m\times 100m)$ outside the users’ area. Note that if Eve is too close to one of the users, it is mainly impossible to guarantee the positive SR for all the users [39]. In this case, other methods, e.g., encryption or friendly jamming, can provide users’ secrecy [11].
@@
The Alice-to-IRS direct path loss factor is ${\aleph_{\text{{AR}}}}\triangleq\boldsymbol{G}_{\text{\text{A}}}+\boldsymbol{G}_{\text{IRS}}-35.9-22\log_{10}(d_{{}_{\text{{AR}}}})$ in dB, where $d_{{}_{\text{{AR}}}}$ is the distance between Alice and the IRS in meters, while $\boldsymbol{G}_{\text{\text{A}}}$ is Alice antenna gain, and $\boldsymbol{G}_{\text{IRS}}$ is the IRS elements’ antenna gain [40]. The path loss factor from the IRS to the users and the Eve is ${\aleph_{\text{R}i}}\triangleq\boldsymbol{G}_{\text{IRS}}-33.05-30\log_{10}(d_{\text{R}i})$ dB, for $i\in\{k,e\}$, $d_{\text{R}i}$ is the distance between the IRS and the users and Eve in meters. The spatial correlation matrix is $\boldsymbol{R}_{Ri}]_{l,\bar{l}}\triangleq\exp{(j\pi(l-\bar{l})\sin\hat{\vartheta}\sin\hat{\aleph})}$ for $i\in\{k,e\}$, where $\hat{\aleph}$ is the elevation angle and $\hat{\vartheta}$ is the azimuth angle [17]. The elements of the Alice-to-IRS channel are generated by $\boldsymbol{G}_{{}_{\text{{AR}}}}]_{a,b}\triangleq\exp({j\pi\left((b-1)\sin\overline{\Theta}_{b}\sin\overline{\vartheta}_{b}+(a-1)\sin(\theta_{m})\sin(\vartheta_{b})\right)})$, where ${\Theta}_{n}\in(0,2\pi)$, ${\vartheta}_{n}\in(0,2\pi)$, and $\overline{\Theta}_{n}\triangleq\pi-\theta_{m}$, and $\overline{\vartheta}_{n}\triangleq\pi+\vartheta_{n}$ [17]. The small-scale fading channel gain $\widehat{\boldsymbol{l}}_{i}$ for $i\in\{k,e\}$ is modeled as a Rician fading channel with K-factor=3=3.
@@
We define the CSI error bounds as $\Omega_{i}\triangleq\delta_{i}\|\widehat{\boldsymbol{l}}_{i}\|_{2},\forall k$, where $\delta_{i}\in[0,1),i\in\{k,e\}$ is the relative amount of CSI uncertainty. When $\delta=0$, Alice can obtain the perfect CSI of the IRS to users/Eve reflected channel. In the FBR case, we set $\tau_{\max}=\tau_{e_{\max}}=10^{-5}$. The packet length $N_{t}$ is defined by the transmission duration $t_{t}$ and the bandwidth $\mathcal{B}$, as previously defined $N_{t}\triangleq\mathcal{B}t_{t}$. Hence, we set the transmission duration as $t_{t}=0.1$ ms, which is suitable for FBR transmission [41]. The choice of $0.1$ ms end-to-end delay ensures having a quasi-static channel during FBR communication [42]. The bandwidth $\mathcal{B}$ is set at $1$ MHz. Unless stated otherwise, the simulation parameters are defined in Table I. We multiply the results by $\log_{2}(e)$ to convert them to bps/Hz. Lastly, for comparison purposes, we compare our Max-Min algorithm against the SSR maximization algorithms. Due to the page limitation, the steps to maximize SSR has been omitted, however, the SSR maximization can be tackled by setting the objective function in problem ($\mathcal{P}1$) and problem ($\mathcal{P}2$) to $\underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\sum_{k}^{\mathcal{K}}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta})$.
@@
Figure 2: Convergence rate of the LBR/FBR Max-Min algorithm under perfect/Imperfect CSI with $N=10$, $\mathcal{K}=6$, $M=16$.
@@
Fig. 3 plots the users’ SR distribution with $N=10$, $\mathcal{K}=6$, and $M=16$ under the Max-Min algorithms. The Max-Min algorithms achieve almost the same SR for all the users when the IRS’s PREs are optimized. However, when the IRS’s PREs are not practically optimized, the algorithms fail to achieve secrecy fairness among the users. The results demonstrate the importance of properly optimizing the IRS’s PREs. The proposed algorithm for the FBR case achieves secrecy and SR fairness among the users. The results demonstrate the robustness of the proposed algorithms to achieve secrecy for all the users, even under the FBR constraints, namely, the transmission duration and the latency.
Figure 3: Users’ SR with $N=10$, $\mathcal{K}=6$, $M=16$.
Figure 4: Users’ SR with with $N=10,\mathcal{K}=6,M=16$.
Fig. 4 portrays the users’ SR under different values of the IRS to the users/Eve’s imperfect CSI. The Max-Min algorithm achieves secrecy fairness in the LBR/FBR cases, even when Eve’s reflected channel has higher uncertainty than the users’ reflected channel. The results show that the proposed algorithm is able to achieve secrecy guarantee even under a high level of uncertainty, which demonstrates the robustness of the proposed algorithm.
Figure 5: Jain’s Index Vs. the uncertainty level $\delta_{k}=\delta_{e}$ with $N=10,\mathcal{K}=5,M=16$.
One way to evaluate the degree of fairness in the proposed algorithms is by using the Jain’s index. Jain’s index is defined as, $\text{Jain's Index}=\frac{\left(\sum_{k=1}^{\mathcal{K}}\mathcal{S}_{\mathcal{K}}(\boldsymbol{w},\boldsymbol{\theta})\right)^{2}}{\mathcal{K}\sum_{k=1}^{\mathcal{K}}(\mathcal{S}_{k}(\boldsymbol{w},\boldsymbol{\theta}))^{2}}$, and it is bounded in $1/\mathcal{K},1]$, where the higher value indicates a better fairness [43]. Fig. 5 shows Jain’s index against the relative amount of CSI uncertainty $\delta$. The Max-Min algorithm achieves almost one in the LBR/FBR cases, while the SSR counterpart has nearly zero. The results demonstrate the Max-Min algorithm’s robustness in achieving secrecy fairness among all the users, even in the FBR environment.
@@
Figure 6: Minimum SR Vs. the uncertainty level $\delta_{k}=\delta_{e}$ with $N=10,\mathcal{K}=5,M=16$.
Figure 7: Minimum SR Vs. the number of users $\mathcal{K}$ with $N=10,M=16$, $\delta_{k}=\delta_{e}=0.02$.
@@ VI Conclusions
In this paper, we proposed a framework to achieve secure communications for all the users under the FBR constraints for IoT setting in URLLC applications aided by IRS. Specifically, through linearization and different nonconvex optimization techniques, we designed computationally efficient algorithm to maximize the minimum SR among all users under both perfect and imperfect CSI from IRS to the users and the eavesdropper. Extensive simulations results showed that the Max-Min algorithm can provide secure communications for all the users under FBR constraints even with only imperfect CSI. Note that if Eve is too close to one user, it is infeasible to guarantee secrecy for all the users and other solutions like cryptography-based or friendly jamming-based methods should be in place. In the future, one can consider a multi-hop scenario with the joint coding over multiple IRSs, similar to the cooperative MIMO setting [44].
@@ Appendix A Proof of theorem 1
First we prove that the sequence $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n+1)})$ is non decreasing for all $k$, i.e. $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n+1)})\geq\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})$ for all $n>0$. To that end, with the aid of the CVX solver, we obtain $\boldsymbol{w}^{(n+1)}$ by solving problem ($\mathcal{P}1.3$). CVX solver is guaranteed to find the optimal solution. Thus, we have $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta})>\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n)},\boldsymbol{\theta})$ for all $k$. Next, from ($\mathcal{P}2$), we obtain $\boldsymbol{\theta}^{(n+1)}$ by solving (33), where we have $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w},\boldsymbol{\theta}^{(n+1)})>\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w},\boldsymbol{\theta}^{(n)})$ as stated before.
@@
$\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n+1)})>\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)}),$ (63)
the optimal sequence $\{(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n+1)})\}$ thus converge to a point $\{({\boldsymbol{w}}^{*},{\boldsymbol{\theta}}^{*})\}$ which is the solution obtained from solving ($\mathcal{P}1.3$) and ($\mathcal{P}1.4$).
[…]
@@
$z_{i}\geq 0,z_{i}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,\forall i.$
@@
$z_{i}\geq 0,z_{i}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,\forall i.$
@@
$z_{i}\geq 0,z_{i}\boldsymbol{T}({\boldsymbol{X}}^{*})=0,\forall i,$
@@ Appendix B Proof of Lemma 1
Let $q$ be a scalar complex variable, and $q^{(n)}$ is the fixed point obtained at iteration $(n)$, then the following inequality holds [18]:
$|q|^{2}\geq q^{*,(n)}q+q^{*}q^{(n)}-q^{*,(n)}q^{(n)}.$ (67)
By replacing $q$ with $({\boldsymbol{l}}_{k}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k}$ we obtain (39). Thus, this concludes the proof. $\blacksquare$
@@ Appendix C Proof of theorem 2
Similar to the proof of Theorem 1, it can be shown the sequence $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n+1)})$ is non-decreasing for all $k$, i.e. $\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n+1)})\geq\mathcal{S}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})$ for all $n>0$, and the converged point is a locally optimal solution for problems. $\blacksquare$.